# Constructions are Revealed in Word Distributions

### v1

@@ Abstract
Construction grammar posits that constructions (form-meaning pairings) are acquired through experience with language (the distributional learning hypothesis). But how much information about constructions does this distribution actually contain? Corpus-based analyses provide some answers, but text alone cannot answer counterfactual questions about what caused a particular word to occur. For that, we need computable models of the distribution over strings—namely, pretrained language models (PLMs). Here we treat a RoBERTa model as a proxy for this distribution and hypothesize that constructions will be revealed within it as patterns of statistical affinity. We support this hypothesis experimentally: many constructions are robustly distinguished, including (i) hard cases where semantically distinct constructions are superficially similar, as well as (ii) schematic constructions, whose “slots” can be filled by abstract word classes. Despite this success, we also provide qualitative evidence that statistical affinity alone may be insufficient to identify all constructions from text. Thus, statistical affinity is likely an important, but partial, signal available to learners.
@@ 1 Introduction
A growing literature explores the use of PLMs as tools for testing usage-based linguistic theories Piantadosi (2024); Weissweiler et al. (2023); Goldberg (2024); Millière (2024). Two characteristics of this literature motivate the current study. First, studies to-date have largely been interested in PLMs as simulations of the learner, whereas we are interested in PLMs as simulations of the distribution from which a learner samples. These perspectives are not mutually incompatible, but they naturally prioritize different types of analyses. PLM-as-learner prioritizes questions about the model’s behavior (studied using e.g., prompting, Zhou et al. 2024; Scivetti et al. 2025) or representational geometry (studied using e.g., probing, Garcia et al. 2021). PLM-as-distribution prioritizes questions about how the probabilities over words are influenced by context, irrespective of the model’s internal computations (e.g., hidden states). CxG studies that take this perspective are less common (cf., Veenboer and Bloem, 2023), and none to our knowledge use causal methods. Second, current evidence on the learnability of constructions by PLMs is mixed, with some studies reporting success Potts (2023); Mahowald (2023); Misra and Mahowald (2024) and others failure Zhou et al. (2024); Bonial and Tayyar Madabushi (2024); Scivetti et al. (2025); Weissweiler et al. (2024).
@@
To address these limitations, we develop intervention methods inspired by collostructional analysis Stefanowitsch and Gries (2003)—corpus methods that measure the statistical affinities that constructions induce between the lexical items they contain. Using a PLM (RoBERTa; Liu et al., 2019) as a simulation model, we show that our method—using only the PLM’s distribution—recovers constructional information across diverse construction types, including in previously reported failure cases Zhou et al. (2024) and across the constructional spectrum from substantive (containing slot(s) that are “fixed” to a specific word) to schematic (containing slot(s) that admit abstract classes of words) constructions. Nonetheless, we argue from a combination of first principles and qualitative evidence that our distributional approach is likely insufficient to infer the full constructicon from data. Our contributions are summarized below:
@@
  • Intervention methods, which reveal constructions as patterns of statistical affinity
  • Resolution of previously reported challenges
  • Generalization of the methods to a wide range of other construction types
@@
  • Qualitative analysis, showing the limits of purely distributional approaches
@@ 2 Background
Prior research on constructions in PLMs has largely used probing or prompting. Probing has been used to study the representation of constructions in both sentence and contextualized word embeddings (Weissweiler et al. 2022; Li et al. 2022; See also Weissweiler et al. 2023 for broader survey). Prompting has been used to elicit acceptability judgements Mahowald (2023), semantic understanding Weissweiler et al. (2024), and constructional similarity judgements Bonial and Tayyar Madabushi (2024). As these methods test the distributional learning hypothesis only indirectly, they may be susceptible to false negatives and positives. Successful prompting typically requires models to have logical, metalinguistic, or instruction-following abilities above and beyond basic representation McCoy et al. (2019, 2023); Basmov et al. (2023), and thus prompting might fail to recover constructions that are, in fact, represented in the model’s distribution. Likewise, probing may either fail to capture representational distinctions that are mismatched to the design of the probe Adi et al. (2017) or successfully capture distinctions that make no causal contribution the model’s behavior Belinkov (2022); Hewitt and Liang (2019).
@@
We are motivated to study the distributional encoding of constructions not only by the methodological considerations above, but by prior studies of challenging cases: constructional properties that seem difficult to infer from distributions. For example, Zhou et al. (2024) report that PLMs fail to distinguish the following three superficially similar but semantically distinct constructions (example adapted from their work):
  • Epistemic Adjective Phrase (EAP):
@@
  • Affective Adjective Phrase (AAP):
@@
  • Causal Excess Construction (CEC):
@@
$[~{}[\text{NP}]~{}[\text{V}]~{}\text{so}~{}[\text{ADJ}]~{}]_{1}~{}\text{that}~{}[\text{S}]_{2}$
where the causal semantics of the three differ: there is no causal relation between $1$ and $2$ in EAP, there is causation from $2\rightarrow 1$ in AAP, and vice versa ($1\rightarrow 2$) in CEC. Zhou et al. probe and prompt LMs (GPT-3.5/4, OpenAI et al. 2024; Llama 2, Touvron et al. 2023) and argue that LMs fail to reliably distinguish the CEC from the EAP and AAP (see Section 4 for details). Relatedly, prior studies of other diverse construction types have suggested that schematic constructions with slots for abstract categories or classes, rather than fixed words, may also be especially difficult (e.g., Weissweiler et al., 2022). In this work, we revisit many of these cases and find that distributions often contain strong signals even for subtle constructional properties.
@@ 3 Methods
To test the hypothesis that constructions are revealed as patterns of statistical affinities, we develop two methods that compare output distributions under input interventions. The first considers global affinity: interaction between a single word and the entire context. The second considers local affinity: pairwise interactions between words.111 All code and data are provided at https://github.com/jsrozner/cxs_are_revealed.
We use the RoBERTa language model Liu et al. (2019) for three reasons: (1) it is an open-source, open-weight model with known training data, (2) it is a less performant model than those used by e.g., Zhou et al. (2024), thus providing a conservative test of the distributional learning hypothesis, and (3) it is bidirectional. Although bidirectionality is implausible for process models of language comprehension Frazier and Fodor (1978); Elman (1990); Tanenhaus et al. (1995); Altmann and Mirković (2009); Smith and Levy (2013), our goal is not to study processing but the underlying distribution, and the constructions we explore here depend on subsequent context. And though human language learners see much less data than PLMs, they also engage in active learning, taking action in the world and in conversational exchanges Frank (2023); thus a human learner may have some ability to sample more strategically from the overall distribution than a PLM trained passively on text. For simplicity, we analyze only single-token words, leaving multi-token generalization of these methods to future work.
@@ 3.1 Global Affinity
Define $\mathcal{P}^{i}_{s\setminus\mathcal{I}}$ to be the probability distribution given by the model for the $i$th position in the masked string, $s\setminus\mathcal{I}$. Then global affinity is simply the probability assigned to a word in the bidirectional context:
$\mathcal{P}^{i}_{s\setminus\{i\}}(w_{i})$
When a word has high global affinity, the context (potentially involving a construction) strongly informs Shannon (1948) the word’s identity. Figure 9 shows per-word global affinities for the sentence, My favorite band is Green Day.
@@ 3.2 Local Affinity
Global affinity alone sheds no light on which parts of context affect the model’s output distribution for a particular word position. This is limiting for the study of constructions because constructions often involve interactions between multiple slots. For example, the NPN construction (e.g., day by day, Jackendoff 2008) introduces an interaction between the pair of nouns: the nouns are mutually constrained to be the same. We quantify such pairwise interactions via local affinity between two positions, $i$ and $j$, defined as the distributional difference at position $i$ in a string $s$ as a function of whether the word at position $j$ was masked:
$a_{i,j}=\text{JSD}(\mathcal{P}^{i}_{s\setminus\{i\}},\mathcal{P}^{i}_{s\setminus\{i,j\}})$
where JSD represents Jensen-Shannon divergence Lin (1991). As we are comparing model outputs under intervention (masking), this formula may recall the interchange intervention formula of Geiger et al. (2022). Computing the affinity between each pair of words in an input of length, $n$, words results in an $n\times n$ affinity matrix, from which constructions’ patterns of affinity can be quantified and visualized (see e.g., Figure 3).
@@ 4 Revisiting a Challenging Case
Figure 2: Affinity plot. Row $i$, Col $j$ is obtained by computing JSD for the $j$th word when the $i$th word is/is not masked. so is more affected by that2 (CEC) than by that1.
Figure 3: UMAP projection (EAP: green, AAP: orange) using 5 pairwise affinities. Separability with SVM (dashed) suggests interaction patterns differ for EAP and AAP.
We first use our methods to address the challenge of distinguishing the CEC from the EAP and AAP, which Zhou et al. (2024) test using probing and prompting. Zhou et al. perform classification by probing sentence (GPT-3.5, LLama-2) and adjective (Llama-2) embeddings. They use prompting (GPT-3.5/4, Llama-2) not for classification, but instead in an experiment that suggests that the models do not understand the causal entailments of the constructions. In this section, we show that the affinity methods not only robustly distinguish the constructions, but that they also identify mislabeled examples in the original dataset (see  4.1, 4.3).
@@ 4.1 Models distinguish the CEC from the EAP and AAP in their output distributions
Whereas the so in the EAP and AAP can be replaced by other adverbial modifiers (e.g., very), so is required for the CEC to be grammatical Kay and Sag (2012); Zhou et al. (2024). If RoBERTa distinguishes the CEC from the EAP and AAP then so should be constrained (have high global affinity) in the CEC but not in the EAP and AAP.
We calculate the global affinity for so in each sentence from the dataset and show that the score distinguishes the CEC: Figure 3 shows that any threshold between 0.6 and 0.9 almost perfectly separates the two classes. Thresholding using global affinity at 0.78 correctly characterizes 272/277 sentences (98.2%). Zhou et al. (2024) separately classify CEC vs EAP/AAP and report percentage accuracies of 79.3, 86.5, 68.5 for CEC vs EAP and 78.8, 86.5, 68.0 for CEC vs AAP using GPT, Llama, and LLama-adj, respectively. Given the strength of our results, we did not reimplement their procedure; it is clear that our untrained approach produces a better result.222 In fact, Zhou et al. upsample the underrepresented EAP and AAP classes. As our threshold correctly categorizes 100% of EAP and AAP examples—miscategorizing only 5 CEC examples— upsampling would improve our accuracy.
Out of 277 examples, 11 originally appeared to be misclassified, but upon review, three of them are mislabeled and three are invalid. For example, “This was so funny that I had to buy another copy and read it to my better half,” was originally labeled AAP. We have reported results and produced plots with labels corrected; details of corrected examples and plots without correction are in Appendix B.2. The clear distinction in the model’s distribution between the CEC and EAP/AAP—which are superficially indistinguishable—provides strong evidence that PLMs do in fact “retrieve and use meanings associated with patterns involving multiple tokens” (cf. Weissweiler et al., 2023).
@@ 4.2 Models capture causal relations in the CEC
We have shown that global affinity challenges a previously reported failure of PLM construction representation. Now we push this finding a step further by using affinity matrices to assess how well PLMs capture the causal semantics of the CEC, as revealed by cases involving multiple clausal complements. Consider “I was so excited that1 I saw you that2 I told my Mom.” Given the requirement of so to license the CEC, we hypothesize that if the model captures the causal semantics of the CEC, then so will have a greater affinity to the causal that2 than the affective one (see Figure 3).
To test this hypothesis in general, we draw CEC instances from Zhou et al. (2024) and insert additional complementizer ($\ldots$) phrases to create a small multi-that dataset of 31 test sentences (see Appendix B.3.1). Across all 31 sentences—even one with five that-phrases—we observe perfect correspondence: so always has the highest affinity with the that at the beginning of the causal excess clause. This result suggests that the distribution both distinguishes the CEC from similar constructions, and also provides signal for the underlying semantics.
@@ 4.3 Affinity patterns distinguish the EAP and AAP
Lastly we consider whether the model distinguishes the EAP and AAP. Zhou et al. (2024) showed that the EAP and AAP can be reasonably distinguished, reporting classifier accuracies of 77.1, 71.7, 84.3 for GPT, Llama, and Llama-adj, respectively. However, given that epistemic and affective adjectives do themselves differ, has the probe recovered a constructional distinction or has it just recovered the adjective class? We cannot answer this question via global affinity: unlike the CEC, which is characterized by the fixed slot constraint for so, no words in the EAP/AAP are fixed, and in some cases when the adjective is masked, both the EAP and AAP are possible (e.g., I was so (happy | certain) that I saw you). We therefore instead compare them by analyzing patterns in the affinity matrix.
First, we align each EAP and AAP sentence so that we can directly compare each example: we identify (subj1, verb1, so, adj, that, subj2, verb2) in each sentence. For example in I was so happy that I saw you, we have (I, was, so, happy, that, I, saw). For each example we extract all pairwise affinities between these seven positions. We test the hypothesis that the constructions have distinct signatures of internal interactions by examining whether the two classes cluster together using UMAP McInnes et al. (2018), a low dimensional projection. A UMAP projection of all 49 ($7\times 7$) dimensions does not separate the two classes; however, this projection considers all affinities equally, even those which may not be related to the construction (see Figure 6, e.g.). To address this, we identify potentially salient differences in the EAP and AAP interaction patterns and produce the UMAP plot in Figure 3 using only the five most substantially different affinities across 26 AAP and 14 EAP examples (admittedly few examples; see Appendix B.4 for full details).
This local affinity method, in addition to suggesting separability using patterns of causal interaction between parts of the input, also identifies three mislabeled examples (see B.4). Zhou et al. reported moderate success using probing to distinguish the EAP and AAP and poor results when using prompting to assess models’ knowledge of entailment relations. In contrast, our results suggest that the EAP and AAP are distinguished in the distribution, even without access to the adjective’s identity.
5 Generalizing to Other Constructions

### v2

@@ Abstract
Construction grammar posits that constructions, or form-meaning pairings, are acquired through experience with language (the distributional learning hypothesis). But how much information about constructions does this distribution actually contain? Corpus-based analyses provide some answers, but text alone cannot answer counterfactual questions about what caused a particular word to occur. This requires computable models of the distribution over strings—namely, pretrained language models (PLMs). Here, we treat a RoBERTa model as a proxy for this distribution and hypothesize that constructions will be revealed within it as patterns of statistical affinity. We support this hypothesis experimentally: many constructions are robustly distinguished, including (i) hard cases where semantically distinct constructions are superficially similar, as well as (ii) schematic constructions, whose “slots” can be filled by abstract word classes. Despite this success, we also provide qualitative evidence that statistical affinity alone may be insufficient to identify all constructions from text. Thus, statistical affinity is likely an important, but partial, signal available to learners. [^1] All code and data are provided at https://github.com/jsrozner/cxs_are_revealed.
@@
Figure 1: In s=s=It is a secret so do not spill the beans”, masking beans (1) gives a constrained distribution, where $\mathcal{P}^{(10)}_{s\setminus\{10\}}(\textit{beans})$ is high, so beans has high global affinity (3). By also masking, e.g., spill (2), we get $\mathcal{P}^{(10)}_{s\setminus\{8,10\}}$, compute JSD, and find the words that constrain beans and thus have high local affinity (4).
@@ 1 Introduction
A growing literature explores the use of PLMs as tools for testing usage-based linguistic theories Weissweiler et al. (2023); Goldberg (2024); Millière (2024); Futrell and Mahowald (2025). Two characteristics of this literature motivate the current study. First, studies to-date have largely been interested in PLMs as simulations of the learner, whereas we are interested in PLMs as simulations of the distribution from which a learner samples. These perspectives are not mutually incompatible, but they naturally prioritize different types of analyses. PLM-as-learner prioritizes questions about the model’s behavior (studied using, e.g., prompting, Zhou et al. 2024; Scivetti et al. 2025) or representational geometry (studied using e.g., probing, Garcia et al. 2021). PLM-as-distribution prioritizes questions about how the probabilities over words are influenced by context, irrespective of the model’s internal computations (e.g., hidden states). CxG studies that take this perspective are less common (cf., Veenboer and Bloem, 2023), and none to our knowledge use causal methods. Second, current evidence on the learnability of constructions by PLMs is mixed, with some studies reporting success Potts (2023); Mahowald (2023); Misra and Mahowald (2024) and others failure Zhou et al. (2024); Bonial and Tayyar Madabushi (2024); Scivetti et al. (2025); Weissweiler et al. (2024).
To address these limitations, we draw inspiration from two areas of research: collostructional analysis Stefanowitsch and Gries (2003) and intervention methods (see, e.g., Feder et al., 2021; Geiger et al., 2022). Collostructional analysis measures the statistical affinities that constructions induce between lexical items in a corpus. Intervention methods systematically alter inputs or hidden states and examine the effects on model behavior. In this study, we extend perturbed masking (Wu et al., 2020; Hoover et al., 2021) to develop affinity methods, which leverage PLMs as computable models of the language distribution, thereby extending the correlational methods of collostructional analysis to counterfactual questions about what causes a particular word to occur.
Our core hypothesis—that constructions will be revealed in the distribution—is partially motivated by idioms, which are loosely defined as semi-fixed multi-word expressions with non-compositional meaning (Nunberg et al. 1994; Croft and Cruse 2004, p. 248-53; Espinal and Mateu 2019). When an idiom is “activated” by the surrounding semantics, any compositional or conventional reading is precluded (Hoffmann, 2022, p. 169), thus constraining the slots of the idiom and motivating our notion of global affinity; see Methods. Some prior work has investigated how PLMs capture the non-compositional aspects of idioms Zeng and Bhat (2021); Socolof et al. (2022); He et al. (2025). Since many constructions exhibit some degree of non-compositionality (Croft and Cruse, 2004, p. 248–253), methods that reveal constraints in the distribution might recover a variety of constructions (Croft and Cruse 2004; Taylor 2004; Wulff 2013).
@@
Using a PLM (RoBERTa; Liu et al., 2019) as a simulation model, we show that affinity methods—using only the PLM’s distribution—recover constructional information across diverse construction types, including in previously reported failure cases (cf. Zhou et al., 2024) and across the constructional spectrum from substantive (containing slot(s) that are “fixed” to a specific word) to schematic (containing slot(s) that admit abstract classes of words) constructions. Nonetheless, we argue from a combination of first principles and qualitative evidence that this distributional approach is likely insufficient to infer the full constructicon from data. In this study, we claim that constructions are revealed in word distributions via affinity methods, and we organize our contributions as follows:
@@
  • Extension of prior work (perturbed masking) as affinity methods that reveal constructions as patterns of statistical interaction (§3)
  • Resolution of previously reported challenges using the methods (§4)
  • Generalization of the methods to a wide range of other construction types (§5, §6)
@@
  • Qualitative analysis to characterize method behavior and inform the limits of purely distributional approaches (§7)
@@ 2 Background
Prior research on constructions in PLMs has largely used probing or prompting. Probing has been used to study the representation of constructions in both sentence and contextualized word embeddings (Weissweiler et al. 2022; Li et al. 2022; see also Weissweiler et al. 2023 for broader survey). Prompting has been used to elicit acceptability judgements Mahowald (2023), semantic understanding Weissweiler et al. (2024), and constructional similarity judgements Bonial and Tayyar Madabushi (2024). As these methods test the distributional learning hypothesis only indirectly, they may be susceptible to false negatives and positives. Successful prompting typically requires models to have logical, metalinguistic, or instruction-following abilities above and beyond basic representation McCoy et al. (2019, 2024); Basmov et al. (2024), and thus prompting might fail to recover constructions that are, in fact, represented in the model’s distribution. Likewise, probing can fail to identify relevant representational distinctions that are mismatched to the design of the probe Adi et al. (2017), or recover distinctions that actually make no causal contribution to the model’s behavior Belinkov (2022); Hewitt and Liang (2019).
@@
We are motivated to study the distributional encoding of constructions not only by the methodological considerations above, buy by prior work on challenging constructions that seem difficult to infer from distributions. For example, Zhou et al. (2024) report that PLMs fail to distinguish the following three superficially similar but semantically distinct constructions:
  • Epistemic Adjective Phrase (EAP)
@@
  • Affective Adjective Phrase (AAP)
@@
  • Causal Excess Construction (CEC)
@@
$\text{NP}]~[\text{V}]~\text{so}~[\text{ADJ}]~]_{1}~\text{that}~[\text{S}]_{2}$
where the causal semantics of the three differ: there is no causal relation between $1$ and $2$ in EAP, there is causation from $2\rightarrow 1$ in AAP, and vice versa ($1\rightarrow 2$) in CEC. Zhou et al. probe and prompt LMs (GPT-3.5/4, OpenAI et al. 2024; Llama2, Touvron et al. 2023) and argue that LMs fail to reliably distinguish the CEC from the EAP and AAP; see Section 4. Relatedly, prior studies of other diverse construction types have suggested that schematic constructions with slots for abstract categories or classes, rather than fixed words, may also be especially difficult (e.g., Weissweiler et al., 2022). In this work, we revisit many of these cases and find that distributions often contain strong signals even for subtle constructional properties.
@@ 3 Methods
To test the hypothesis that constructions are revealed as patterns of statistical affinities, we extend perturbed masking Wu et al. (2020); Hoover et al. (2021), developing two approaches that compare output distributions under input interventions. The first is global affinity: interaction between a single word and the entire context. The second is local affinity: pairwise interactions between words.
We use the RoBERTa language model Liu et al. (2019) for three reasons: (1) it is an open-source, open-weight model with known training data and a pure language modeling objective (e.g., no instruction tuning), (2) it is a less performant model than those used by e.g., Zhou et al. (2024), thus providing a conservative test of the distributional learning hypothesis, and (3) it is bidirectional. Although bidirectionality is implausible for process models of language comprehension Frazier and Fodor (1978); Elman (1990); Tanenhaus et al. (1995); Altmann and Mirković (2009); Smith and Levy (2013), our goal is not to study processing but the underlying distribution, and the constructions we explore here depend on subsequent context. And though human language learners see much less data than PLMs, they also engage in active learning, taking action in the world and in conversational exchanges Frank (2023); thus a human learner may have some ability to sample more strategically from the overall distribution than a PLM trained passively on text. For simplicity, we analyze only single-token words, leaving multi-token generalization of these methods to future work. As RoBERTa has a relatively large vocabulary (50k), we find that this does not pose a substantial limitation for our study.
@@ 3.1 Global Affinity
Define $\mathcal{P}^{(i)}_{s\setminus\mathcal{I}}$ to be the probability distribution given by the model for the $i$th position in the masked string, $s\setminus\mathcal{I}$ (note that $i\in\mathcal{I}$). Then global affinity is simply the probability assigned to the original word in the bidirectional context:
$\mathcal{P}^{(i)}_{s\setminus\{i\}}(w_{i})$
When a word has high global affinity, the context—potentially involving a construction—strongly informs Shannon (1948) the word’s identity. Figure 10 shows per-word global affinities for the sentence, My favorite band is Green Day.
@@ 3.2 Local Affinity
Global affinity alone sheds no light on which parts of context affect the model’s output distribution for a particular word position. This is limiting for the study of constructions because constructions often involve interactions between multiple slots. For example, the NPN construction (e.g., day by day, Jackendoff 2008) introduces an interaction between the pair of nouns: the nouns are mutually constrained to be the same. We quantify such pairwise interactions via local affinity between two positions, $i$ and $j$, defined as the distributional difference at position $j$ in a string $s$ as a function of whether the word at position $i$ was masked:
$a_{i,j}=\text{JSD}(\mathcal{P}^{(j)}_{s\setminus\{j\}},\mathcal{P}^{(j)}_{s\setminus\{i,j\}})$
where JSD represents Jensen-Shannon divergence Lin (1991). Computing the affinity between each pair of words in an input of length, $n$, words results in an $n\times n$ affinity matrix, from which constructions’ patterns of affinity can be quantified and visualized (see e.g., Figure 4).
@@ 4 Revisiting a Challenging Case
Figure 3: Local Affinity plot. A column shows how much the distribution for that word is affected by other words in the context. so is more affected by that2 (CEC) than by that1.
Figure 4: UMAP projection (EAP: orange, AAP: blue) using 5 pairwise affinities. Separability with SVM (dashed) suggests interaction patterns differ for EAP and AAP.
We first use our methods to address the challenge of distinguishing the CEC from the EAP and AAP, which Zhou et al. (2024) test using probing and prompting. Zhou et al. perform classification by probing sentence (GPT-3.5, LLama2) and adjective (Llama2) embeddings. They use prompting (GPT-3.5/4, Llama2) not for classification, but instead in an experiment that suggests that the models do not understand the causal entailments of the constructions. In this section, we show that the affinity methods not only robustly distinguish the constructions, but that they also identify mislabeled examples in the original dataset (see  §4.1, §4.3).
@@ 4.1 Models distinguish the CEC from the EAP and AAP in their output distributions
Whereas the so in the EAP and AAP can be replaced by other adverbial modifiers (e.g., very), so is required for the CEC to be grammatical Kay and Sag (2012); Zhou et al. (2024). If RoBERTa distinguishes the CEC from the EAP and AAP, then so should be constrained (have high global affinity) in the CEC but not in the EAP and AAP.
We calculate the global affinity for so in each sentence and observe that the score distinguishes the CEC: thresholding global affinity at 0.78 correctly characterizes 272/277 sentences (98.2%). In fact, Figure 4 shows that there is a wide margin, since any threshold between 0.6 and 0.9 achieves similar separation. Zhou et al. (2024) separately classify CEC vs. EAP/AAP and report percentage accuracies of 79.3, 86.5, 68.5 for CEC vs. EAP and 78.8, 86.5, 68.0 for CEC vs. AAP using GPT, Llama, and LLama-adj, respectively. Given the strength of our results, we did not reimplement their procedure; it is clear that our untrained approach produces a better result. [^2] In fact, Zhou et al. upsample the underrepresented EAP and AAP classes. As our threshold correctly categorizes 100% of EAP and AAP examples—miscategorizing only 5 CEC examples— upsampling would improve our accuracy.
Out of 277 examples, 11 originally appeared to be misclassified using the 0.78 threshold, but upon review, three were mislabeled and three were invalid. For example, “This was so funny that I had to buy another copy and read it to my better half,” was originally labeled AAP. We report results and plots with labels corrected; we provide details of corrected examples in Appendix B.2. The clear distinction in the model’s distribution between the CEC and EAP/AAP—which are superficially indistinguishable—provides strong evidence that PLMs do in fact “retrieve and use meanings associated with patterns involving multiple tokens” (cf. Weissweiler et al., 2023).
@@ 4.2 Models capture causal relations in the CEC
We have shown that global affinity challenges a previously reported failure of PLM construction representation. Now we push this finding a step further by using affinity matrices to assess how well PLMs capture the causal semantics of the CEC, as revealed by cases involving multiple clausal complements. Consider I was so excited that1 I saw you that2 I told my Mom. Given the requirement of so to license the CEC, we hypothesize that if the model captures the causal semantics of the CEC, then so will have a greater affinity to the causal that2 than to the affective one (see Figure 4).
[…]
@@ 4.3 Affinity patterns distinguish the EAP and AAP
Lastly we consider whether the model distinguishes the EAP and AAP. Zhou et al. (2024) showed that the EAP and AAP can be reasonably distinguished, reporting classifier accuracies of 77.1, 71.7, 84.3 for GPT, Llama, and Llama-adj, respectively. However, given that epistemic and affective adjectives do themselves differ, has the probe recovered a constructional distinction or has it just recovered the adjective class? We cannot answer this question via global affinity: unlike the CEC, which is characterized by the fixed slot constraint for so, no words in the EAP/AAP are fixed, and in some cases when the adjective is masked, both the EAP and AAP are possible (e.g., I was so (happy | certain) that I saw you). We therefore instead compare them by analyzing patterns in the local affinity matrix.
First, we align examples by identifying the following parts common to all inputs: <subj1, verb1, so, adj, that, subj2, verb2>. For example in I was so happy that I saw you, we have <I, was, so, happy, that, I, saw>. For each example we extract all pairwise affinities between these seven positions. We test the hypothesis that the constructions have distinct signatures of internal interactions by examining whether the two classes cluster together using UMAP McInnes et al. (2018), a low dimensional projection. A UMAP projection of all 49 ($7\times 7$) dimensions does not separate the two classes; however, this projection considers all affinities equally, even those which may not be related to the construction (see e.g., Figure 7). To address this, we identify potentially salient differences in the interaction patterns and produce the UMAP plot in Figure 4 using only the five most substantially different affinities across 26 AAP and 14 EAP examples (admittedly few examples; see Appendix B.4 for details).
This method, though imperfect, suggests separability using patterns of interaction between parts of the input. (See B.4 for further discussion of separability.) Zhou et al. were moderately successful in distinguishing the EAP and AAP using probing and unsuccessful with prompting. Our results suggest that the distribution may be able to distinguish EAP and AAP examples, even without access to the adjective’s identity.
5 Generalizing to Other Substantive Constructions